OpenStudy (anonymous):

If Y=sin (5x)^-1,then dy/dx=??

OpenStudy (tkhunny):

Is that \(\sin^{-1}(5x)\;or\;\dfrac{1}{\sin(5x)}\)?

OpenStudy (tkhunny):

...or maybe \(\sin\left(\dfrac{1}{5x}\right)\)?

OpenStudy (anonymous):

first one

OpenStudy (tkhunny):

Find \(\dfrac{d}{dx}\sin^{-1}(x)\)